# NeuroBOLT: Resting-state EEG-to-fMRI Synthesis with Multi-dimensional Feature Mapping

### v2

@@ 1 Introduction
Functional magnetic resonance imaging (fMRI) and electroencephalography (EEG) are the most commonly utilized non-invasive neuroimaging techniques, providing crucial insights into brain functionality. These two modalities offer distinct advantages and limitations that can complement one another when combined [45, 13]. fMRI offers high spatial resolution imaging of whole-brain activity by measuring blood-oxygen-level-dependent (BOLD) signal changes, which facilitates probing regional and network-level function. However, fMRI suffers from low temporal resolution and hemodynamic blurring, which limits its ability to capture the accurate timing of rapid neuronal activity dynamics. Additionally, its high cost, non-portability, and incompatibility with metal implants further restrict the utility of MRI in certain contexts. Conversely, EEG stands out as a low-cost, portable neuroimaging modality with high temporal resolution [10, 30, 61]. However, EEG faces limitations due to volume conduction and from the superficial location of electrodes (on the scalp), making it difficult for EEG to accurately infer the origins of the measured electrical potentials. In this context, the high spatial resolution of fMRI becomes indispensable, especially for imaging deep brain regions.
@@
One approach for investigating the relationship between EEG and fMRI signals is by analyzing data collected simultaneously from both modalities. Such studies have demonstrated correlations between fMRI data and various features of EEG signals, such as the power in certain frequency bands [27, 7, 28], which underscore the potential of EEG to inform fMRI features. However, factors such as the disparate biophysical origins of the two signals, and differences in their spatial and temporal resolutions, can limit the accuracy, interpretation, and consistency of correlation-based EEG-fMRI studies. These challenges are exacerbated in conditions during which a participant is resting passively (i.e., resting state), which is characterized by significant noise and randomness. Consequently, the mechanisms linking neuronal activity to BOLD signals remain only partially understood, posing challenges in mapping EEG to fMRI.
@@
With recent progress in cross-modality synthesis and brain decoding techniques using deep neural networks [8, 9, 16], EEG-to-fMRI synthesis has emerged as a promising yet largely untapped research area [4]. Early pioneering work by [39, 43] employed ridge regression on EEG temporal-spectral features to reconstruct fMRI signals from the visual cortex and sub-cortical regions. More recent studies have used deep neural networks for EEG-fMRI translation [33, 34, 35, 5, 3, 25, 31], and those of [25] and [31] developed sequence-to-sequence (Seq-to-Seq) models to reconstruct fMRI time series of deep brain regions from EEG. However, these efforts have primarily focused on task datasets, including cued eye opening/closing [25, 31, 34, 35, 54, 3], leaving the (fully eyes-closed) resting state condition largely unexplored. In addition, prior studies rely on a subject-specific approach, wherein models are solely trained and tested on different sections of the same individual’s scans. Further, only a small number of brain regions have been examined in current studies, leaving as an open question the predictive power of EEG for fMRI signals across broader areas (see Appendix B for further discussion of related work). These limitations underscore the need for developing more generalizable models for EEG-fMRI translation that are accompanied by more comprehensive evaluation.
@@
In this paper, we present NeuroBOLT: a multi-dimensional transformer-based EEG encoding framework for Neural-to-BOLD fMRI Translation (Figure 2). NeuroBOLT is designed to capture and integrate multi-dimensional representations across temporal, spatial, and spectral domains, offering a generalizable approach for translating raw EEG waves to BOLD fMRI time series from regions of interest (ROIs). Drawing inspiration from vision transformer (ViT) [15] and its adaptation for multi-channel biosignals [59, 22], our framework transforms raw EEG data into unified "EEG patches", enabling flexible and effective EEG data encoding. These EEG patches are then passed into two parallel sub-modules: (1) spatiotemporal and (2) spectral representation learning modules. The spatiotemporal module is propelled by a recently proposed EEG foundation model (LaBraM) [22], which has been trained to learn effective representations on about 2,500 hours of various types of EEG signals. Moreover, as previous studies have emphasized the importance of leveraging spectral features in EEG representation learning [59, 57, 60], EEG-fMRI correlations [27, 14], and EEG-to-fMRI synthesis [31, 39, 25], the spectral module incorporates multi-scale encoding on EEG spectrogram patches. Specifically, instead of employing a fixed window size for the Short-Time Fourier Transform (STFT) as commonly used [59, 58, 24, 11], we incorporate spectral features from windows of varying scales. This approach retains the advantages of high temporal resolution from smaller windows and high frequency resolution from larger windows, thereby enhancing the spectral analysis. The output embeddings from the above two encoding modules are then integrated, allowing NeuroBOLT to capture the complexity of neural dynamics and learn the projection from neural to BOLD signals. The key contributions of this work are summarized as follows:
@@ 2.1 Task Formulation
While it is well-established that fMRI signals are coupled to neuronal activity, the details of this process are only partially understood and still under debate [19]. In practice, most studies model the relationship between neuronal activity and fMRI by convolving the assumed neural activity with the hemodynamic response function (HRF), which reaches its peak at about 6 seconds and gradually returns to baseline over the next 12-15 seconds. However, the HRF is not uniform, varying significantly across different brain regions and between individuals. Moreover, although EEG can detect the electrical activity of neural populations at millisecond timescales, it suffers from volume conduction, making it challenging to extract its neural origins. This spatial ambiguity also complicates the interpretation of EEG and its relation to fMRI data.
@@
Functional MRI data are conventionally represented in four dimensions, and can be expressed as a temporal sequence $S=\{V_{1},...,V_{K}\}$ with K observations, where each observation is a volume $V\in\mathbb{R}^{x\times y\times\ z}$ with spatial dimensions $x,y,$ and $z$. This high-dimensional fMRI signal is often summarized by representing it as a set of brain areas (i.e., $P$ parcels), each with a corresponding fMRI signal that is averaged over its constituent voxels $v_{x,y,z}\subset V$. Here, we employ such a functional parcellation, Dictionaries of Functional Modes (DiFuMo) [12], which was learned across millions of fMRI volumes (see details in [12, 50]). We utilize the DiFuMo parcellation with $P=64$ regions, and the resulting parcellated fMRI data $Y\in\mathbb{R}^{P\times K}$ contains the signals from each of these regions.
@@ 2.2 Model Architecture
In this section, we introduce our model: NeuroBOLT, a general architecture for translating EEG signals to fMRI. Our model is designed to accommodate input EEG signals with an arbitrary number of channels. As shown in Figure 2, we leverage the pre-trained EEG foundation model, LaBraM (checkpoint version: LaBraM-base) [22], and finetune on our dataset to obtain spatiotemporal representations of EEG signals. Additionally, we propose multi-scale spectral feature fusion to obtain comprehensive spectral representations. These two modules learn complementary attributes of EEG data. Finally, we integrate the spatiotemporal and multi-scale spectral representations and feed them into a regression head for fMRI prediction.
@@
We formulate the multi-channel EEG signals as $X\in\mathbb{R}^{C\times T}$, where $C$ represents the number of EEG electrode channels and $T$ denotes the time length of the input EEG. To obtain the spatiotemporal representation of a given set of EEG signals, we leverage an operation from LaBraM [22], which segments the EEG signals into patches. Assuming the time window length (patch size) for tokenization of EEG signal is $w$ and the stride is $s$, $X$ can be segmented into $\left\lfloor\frac{T-w}{s}\right\rfloor+1$ segments, with each segment denoted as $\mathbf{x}\in\mathbb{R}^{C\times w}$. In this work, we use a window of length $w$ and a stride $s=w$ to segment each EEG channel into patches, obtaining $\mathbf{x}=\{x_{c_{j},k}\in\mathbb{R}^{w}\mid j=1,2,\dots,C,k=1,2,\dots,\left\lfloor\frac{T}{w}\right\rfloor\}$. The total number of patches is $C\times\left\lfloor\frac{T}{w}\right\rfloor$. These patches are then passed forward to a temporal encoder [22] to obtain the patch embeddings, denoted as:
@@
The segment embedding $e_{seg}$ is directly fed into the Transformer encoder [52] to obtain the output embeddings. We then apply average pooling to these embeddings to obtain the spatial and temporal representation $r_{st}\in\mathbb{R}^{d_{st}}$, where $d_{st}$ denotes the dimensionality.
@@
Compared to RGB images, EEG signals present several challenges, such as a low signal-to-noise ratio, apparent stochasticity, nonstationarity, and nonlinear characteristics, making the reconstruction of the original signals difficult [41]. Previous research indicates that the frequency spectrum of EEG signals is crucial for understanding the brain’s neurophysiological activities [59, 55]. Therefore, in our work, we utilize the Short-Time Fourier Transform (STFT) to achieve a spectral representation of EEG signals. Unlike most recent state-of-the-art methods, we design a multi-scale spectral approach that captures both coarse and fine representations in the temporal and frequency domains [56, 36]. Details of rationale can be found in Appendix C).
@@
For each channel, we apply spatial embedding, similar to the spatiotemporal representation section, to obtain the final embedding with dimensions $e_{sp}\in\mathbb{R}^{C\times n\times d}$, where $C$, $n$ and $d$ represent the channel number, window embedding dimension, and frequency embedding dimension, respectively. The original Transformer is known to have quadratic complexity in both time and space, and EEG signals typically contain tens of channels. To learn these embeddings with lower complexity [59], we feed spectrum embeddings $e_{sp}$ into a linear Transformer Encoder [53, 23]. Let the input embedding $e_{sp}\in\mathbb{R}^{N\times d}$, where $N=C\times n$ represents the number of tokens. The self-attention operation in the linear Transformer Encoder can be formulated as follows:
@@
Here $\textbf{W}^{Q}$, $\textbf{W}^{K}$, $\textbf{W}^{V}$ $\in\mathbb{R}^{d\times k}$ are the query, key and value matrices. The self-attention module uses a rank-$D$ approximation for the softmax attention $(N\times N)$ by reduced-rank parameter matrices $\textbf{E}^{\top}\in\mathbb{R}^{N\times D}$, $\textbf{F}\in\mathbb{R}^{D\times N}$ (where $D\ll N$). Main components of this module include one linear self-attention layer and one fully connected network. Layer normalization before each component [2], residual connection after each component [18], and dropout right after the self-attention to enable stable training [49, 59]. The output is fed into average pooling along the token dimension to obtain the final spectral representation $r_{sp}\in\mathbb{R}^{d_{sp}}$, where $r_{sp}$ has the same dimension as the spatiotemporal representation $r_{st}$.
@@
The hidden embeddings from the above two modules are then summed up and fed into a regression head, which consists of Gaussian Error Linear Unit (GELU) [20] followed by a single linear layer, to make the final prediction of fMRI in the $p^{th}$ ROI at time $t$.
@@ 3.2 Experimental Setup
We employ DiFuMo with $n=64$ dimensions [12] (see Section 2.1) to extract measured BOLD signals within specific ROIs. We focus on seven ROIs that span a range of spatial locations and functional roles, namely: primary sensory regions (cuneus and Heschl’s gyrus), high-level cognitive areas (anterior precuneus, anterior and middle frontal gyri), subcortical regions (putamen and thalamus), and the global (whole-brain average) signal. From these ROI time series, we regress out motion confounds, low-pass filter the signals below 0.15Hz, and use the 95th-percentile of the absolute amplitude to normalize the demeaned ROI time courses. To prepare the EEG data for input, we first exclude the ECG, EOG, and EMG channels (remaining 26 channels), and resample the EEG to 200 Hz to enhance computational efficiency while preserving meaningful frequency components (which are typically below 100 Hz). To predict the fMRI ROI signal, we extract EEG windows of 16 seconds before each fMRI data point. This window length is selected to encompass the peak and most of the variation in the hemodynamic response function. Additionally, this duration aligns with the maximum temporal encoding window length of the pre-trained LaBraM model [22], ensuring optimal fine-tuning. The normalization of the EEG also strictly follows [22], where EEG data values are divided by 100 so that the resulting amplitude falls primarily between -1 to 1, since a typical EEG amplitude range is from -100 $\mu V$ to 100 $\mu V$.
@@
The baseline models include state-of-the-art EEG encoding models from [59] and [22], as well as two EEG-to-fMRI translation baselines [25, 31]. Since the original EEG-to-fMRI baselines are sequence-to-sequence models, we adapted them by adding a final projection layer to their encoder to enable sequence-to-one prediction. Additional details on baseline implementations are provided in the Appendix E.2. We employed the following two metrics for comparison: 1) Pearson correlation coefficient (R), which measures the strength and direction of the linear relationship between prediction and ground truth; 2) Mean squared error (MSE), which measures the average squared differences between prediction and ground truth across samples. The main text highlights the correlation coefficient as the primary evaluation metric, while MSE results are provided in the Appendix Table 6.
@@
For the spatiotemporal module, we initialize the model by loading the pretrained weights from LaBraM-base [22], with a token length = 200, i.e., 1 second, with no overlap. Please refer to [22] for more details of EEG pretraining. For the multi-scale spectral module, we set the smallest scale size $l_{0}=100$, i.e., 0.5 seconds without overlap. Experiments are conducted on a single RTX A5000 GPU using Python 3.9.12, Pytorch 2.0.1, and CUDA 11.7. To ensure consistency during model training, we set a fixed seed across all experiments. The batch sizes are set at 16 and 64 for intra-subject and inter-subject analyses, respectively. AdamW is utilized as our optimizer, and MSE as our training objective. The initial learning rate is set at 3e-4 with a weight decay of 0.05, and a minimal learning rate of 1e-6. For subject-specific prediction, where training and testing occur on the same scan, we split the scan in an 8:1:1 ratio for training, validation, and testing, i.e., training on the first 80% of the data and testing on the last 10%. Given that the fMRI signal exhibits significant autocorrelation, typically extending from about -10 to 10 seconds [6], we implement gaps of 20 seconds between the training and validation sets, as well as between the validation and testing sets, to prevent data leakage. For the inter-subject analysis in resting-state data, we randomly divided the datasets into training/validation/testing sets by approximately 3:1:1 (18 scans : 5 scans : 6 scans). For the task data, we split the scans by 9 scans : 3 scans : 4 scans. Since data from the same individual might have shared representations, the two scans from the same individual are ensured to be in the same set (either training, validation, or testing set) to avoid possible data leakage. All models are optimized on the training set and evaluated on the test set, with the best model and hyperparameters selected based on the validation set.

### v3

@@ 1 Introduction
Functional magnetic resonance imaging (fMRI) and electroencephalography (EEG) are the most commonly utilized non-invasive neuroimaging techniques, providing crucial insights into brain functionality. These two modalities offer distinct advantages and limitations that can complement one another when combined Ritter and Villringer (2006); Daunizeau et al. (2010). fMRI offers high spatial resolution imaging of whole-brain activity by measuring blood-oxygen-level-dependent (BOLD) signal changes, which facilitates probing regional and network-level function. However, fMRI suffers from low temporal resolution and hemodynamic blurring, which limits its ability to capture the accurate timing of rapid neuronal activity dynamics. Additionally, its high cost, non-portability, and incompatibility with metal implants further restrict the utility of MRI in certain contexts. Conversely, EEG stands out as a low-cost, portable neuroimaging modality with high temporal resolution Cohen (2017); Li et al. (2020); Zhang et al. (2021). However, EEG faces limitations due to volume conduction and from the superficial location of electrodes (on the scalp), making it difficult for EEG to accurately infer the origins of the measured electrical potentials. In this context, the high spatial resolution of fMRI becomes indispensable, especially for imaging deep brain regions.
@@
One approach for investigating the relationship between EEG and fMRI signals is by analyzing data collected simultaneously from both modalities. Such studies have demonstrated correlations between fMRI data and various features of EEG signals, such as the power in certain frequency bands Laufs et al. (2003); Chang et al. (2013); Laufs et al. (2006), which underscore the potential of EEG to inform fMRI features. However, factors such as the disparate biophysical origins of the two signals, and differences in their spatial and temporal resolutions, can limit the accuracy, interpretation, and consistency of correlation-based EEG-fMRI studies. These challenges are exacerbated in conditions during which a participant is resting passively (i.e., resting state), which is characterized by significant noise and randomness. Consequently, the mechanisms linking neuronal activity to BOLD signals remain only partially understood, posing challenges in mapping EEG to fMRI.
@@
With recent progress in cross-modality synthesis and brain decoding techniques using deep neural networks Chen et al. (2023, 2024); Duan et al. (2024), EEG-to-fMRI synthesis has emerged as a promising yet largely untapped research area Calhas and Henriques (2020). Early pioneering work by (Meir-Hasson et al., 2014; Or-Borichev et al., 2023) employed ridge regression on EEG temporal-spectral features to reconstruct fMRI signals from the visual cortex and sub-cortical regions. More recent studies have used deep neural networks for EEG-fMRI translation Liu and Sajda (2019, 2023a, 2023b); Calhas and Henriques (2022); Bricman and Borst (2021); Kovalev et al. (2022); Li et al. (2024), and those of Kovalev et al. (2022) and Li et al. (2024) developed sequence-to-sequence (Seq-to-Seq) models to reconstruct fMRI time series of deep brain regions from EEG. However, these efforts have primarily focused on task datasets, including cued eye opening/closing Kovalev et al. (2022); Li et al. (2024); Liu and Sajda (2023a, b); Wei et al. (2020); Bricman and Borst (2021), leaving the (fully eyes-closed) resting state condition largely unexplored. In addition, prior studies rely on a subject-specific approach, wherein models are solely trained and tested on different sections of the same individual’s scans. Further, only a small number of brain regions have been examined in current studies, leaving as an open question the predictive power of EEG for fMRI signals across broader areas (see Appendix B for further discussion of related work). These limitations underscore the need for developing more generalizable models for EEG-fMRI translation that are accompanied by more comprehensive evaluation.
@@
In this paper, we present NeuroBOLT: a multi-dimensional transformer-based EEG encoding framework for Neural-to-BOLD fMRI Translation (Figure 2). NeuroBOLT is designed to capture and integrate multi-dimensional representations across temporal, spatial, and spectral domains, offering a generalizable approach for translating raw EEG waves to BOLD fMRI time series from regions of interest (ROIs). Drawing inspiration from vision transformer (ViT) Dosovitskiy et al. (2020) and its adaptation for multi-channel biosignals Yang et al. (2023); Jiang et al. (2024), our framework transforms raw EEG data into unified "EEG patches", enabling flexible and effective EEG data encoding. These EEG patches are then passed into two parallel sub-modules: (1) spatiotemporal and (2) spectral representation learning modules. The spatiotemporal module is propelled by a recently proposed EEG foundation model (LaBraM) Jiang et al. (2024), which has been trained to learn effective representations on about 2,500 hours of various types of EEG signals. Moreover, as previous studies have emphasized the importance of leveraging spectral features in EEG representation learning Yang et al. (2023, 2021); Yi et al. (2024), EEG-fMRI correlations Laufs et al. (2003); de Munck et al. (2009), and EEG-to-fMRI synthesis Li et al. (2024); Meir-Hasson et al. (2014); Kovalev et al. (2022), the spectral module incorporates multi-scale encoding on EEG spectrogram patches. Specifically, instead of employing a fixed window size for the Short-Time Fourier Transform (STFT) as commonly used Yang et al. (2023, 2022); Kim et al. (2020); Cui et al. (2020), we incorporate spectral features from windows of varying scales. This approach retains the advantages of high temporal resolution from smaller windows and high frequency resolution from larger windows, thereby enhancing the spectral analysis. The output embeddings from the above two encoding modules are then integrated, allowing NeuroBOLT to capture the complexity of neural dynamics and learn the projection from neural to BOLD signals. The key contributions of this work are summarized as follows:
@@ 2.1 Task Formulation
While it is well-established that fMRI signals are coupled to neuronal activity, the details of this process are only partially understood and still under debate Heeger and Ress (2002). In practice, most studies model the relationship between neuronal activity and fMRI by convolving the assumed neural activity with the hemodynamic response function (HRF), which reaches its peak at about 6 seconds and gradually returns to baseline over the next 12-15 seconds. However, the HRF is not uniform, varying significantly across different brain regions and between individuals. Moreover, although EEG can detect the electrical activity of neural populations at millisecond timescales, it suffers from volume conduction, making it challenging to extract its neural origins. This spatial ambiguity also complicates the interpretation of EEG and its relation to fMRI data.
@@
Functional MRI data are conventionally represented in four dimensions, and can be expressed as a temporal sequence $S=\{V_{1},...,V_{K}\}$ with K observations, where each observation is a volume $V\in\mathbb{R}^{x\times y\times\ z}$ with spatial dimensions $x,y,$ and $z$. This high-dimensional fMRI signal is often summarized by representing it as a set of brain areas (i.e., $P$ parcels), each with a corresponding fMRI signal that is averaged over its constituent voxels $v_{x,y,z}\subset V$. Here, we employ such a functional parcellation, Dictionaries of Functional Modes (DiFuMo) Dadi et al. (2020), which was learned across millions of fMRI volumes (see details in Dadi et al. (2020); Thomas et al. (2022)). We utilize the DiFuMo parcellation with $P=64$ regions, and the resulting parcellated fMRI data $Y\in\mathbb{R}^{P\times K}$ contains the signals from each of these regions.
@@ 2.2 Model Architecture
In this section, we introduce our model: NeuroBOLT, a general architecture for translating EEG signals to fMRI. Our model is designed to accommodate input EEG signals with an arbitrary number of channels. As shown in Figure 2, we leverage the pre-trained EEG foundation model, LaBraM (checkpoint version: LaBraM-base) Jiang et al. (2024), and finetune on our dataset to obtain spatiotemporal representations of EEG signals. Additionally, we propose multi-scale spectral feature fusion to obtain comprehensive spectral representations. These two modules learn complementary attributes of EEG data. Finally, we integrate the spatiotemporal and multi-scale spectral representations and feed them into a regression head for fMRI prediction.
@@
We formulate the multi-channel EEG signals as $X\in\mathbb{R}^{C\times T}$, where $C$ represents the number of EEG electrode channels and $T$ denotes the time length of the input EEG. To obtain the spatiotemporal representation of a given set of EEG signals, we leverage an operation from LaBraM Jiang et al. (2024), which segments the EEG signals into patches. Assuming the time window length (patch size) for tokenization of EEG signal is $w$ and the stride is $s$, $X$ can be segmented into $\left\lfloor\frac{T-w}{s}\right\rfloor+1$ segments, with each segment denoted as $\mathbf{x}\in\mathbb{R}^{C\times w}$. In this work, we use a window of length $w$ and a stride $s=w$ to segment each EEG channel into patches, obtaining $\mathbf{x}=\{x_{c_{j},k}\in\mathbb{R}^{w}\mid j=1,2,\dots,C,k=1,2,\dots,\left\lfloor\frac{T}{w}\right\rfloor\}$. The total number of patches is $C\times\left\lfloor\frac{T}{w}\right\rfloor$. These patches are then passed forward to a temporal encoder Jiang et al. (2024) to obtain the patch embeddings, denoted as:
@@
The segment embedding $e_{seg}$ is directly fed into the Transformer encoder Vaswani et al. (2017) to obtain the output embeddings. We then apply average pooling to these embeddings to obtain the spatial and temporal representation $r_{st}\in\mathbb{R}^{d_{st}}$, where $d_{st}$ denotes the dimensionality.
@@
Compared to RGB images, EEG signals present several challenges, such as a low signal-to-noise ratio, apparent stochasticity, nonstationarity, and nonlinear characteristics, making the reconstruction of the original signals difficult Moss et al. (2004). Previous research indicates that the frequency spectrum of EEG signals is crucial for understanding the brain’s neurophysiological activities Yang et al. (2023); Wu et al. (2022a). Therefore, in our work, we utilize the Short-Time Fourier Transform (STFT) to achieve a spectral representation of EEG signals. Unlike most recent state-of-the-art methods, we design a multi-scale spectral approach that captures both coarse and fine representations in the temporal and frequency domains Wu et al. (2022b); Lou et al. (2024). Details of rationale can be found in Appendix C).
@@
For each channel, we apply spatial embedding, similar to the spatiotemporal representation section, to obtain the final embedding with dimensions $e_{sp}\in\mathbb{R}^{C\times n\times d}$, where $C$, $n$ and $d$ represent the channel number, window embedding dimension, and frequency embedding dimension, respectively. The original Transformer is known to have quadratic complexity in both time and space, and EEG signals typically contain tens of channels. To learn these embeddings with lower complexity Yang et al. (2023), we feed spectrum embeddings $e_{sp}$ into a linear Transformer Encoder Wang et al. (2020); Katharopoulos et al. (2020). Let the input embedding $e_{sp}\in\mathbb{R}^{N\times d}$, where $N=C\times n$ represents the number of tokens. The self-attention operation in the linear Transformer Encoder can be formulated as follows:
@@
Here $\textbf{W}^{Q}$, $\textbf{W}^{K}$, $\textbf{W}^{V}$ $\in\mathbb{R}^{d\times k}$ are the query, key and value matrices. The self-attention module uses a rank-$D$ approximation for the softmax attention $(N\times N)$ by reduced-rank parameter matrices $\textbf{E}^{\top}\in\mathbb{R}^{N\times D}$, $\textbf{F}\in\mathbb{R}^{D\times N}$ (where $D\ll N$). Main components of this module include one linear self-attention layer and one fully connected network. Layer normalization before each component Ba (2016), residual connection after each component He et al. (2016), and dropout right after the self-attention to enable stable training Srivastava et al. (2014); Yang et al. (2023). The output is fed into average pooling along the token dimension to obtain the final spectral representation $r_{sp}\in\mathbb{R}^{d_{sp}}$, where $r_{sp}$ has the same dimension as the spatiotemporal representation $r_{st}$.
@@
The hidden embeddings from the above two modules are then summed up and fed into a regression head, which consists of Gaussian Error Linear Unit (GELU) Hendrycks and Gimpel (2016) followed by a single linear layer, to make the final prediction of fMRI in the $p^{th}$ ROI at time $t$.
@@ 3.2 Experimental Setup
We employ DiFuMo with $n=64$ dimensions Dadi et al. (2020) (see Section 2.1) to extract measured BOLD signals within specific ROIs. We focus on seven ROIs that span a range of spatial locations and functional roles, namely: primary sensory regions (cuneus and Heschl’s gyrus), high-level cognitive areas (anterior precuneus, anterior and middle frontal gyri), subcortical regions (putamen and thalamus), and the global (whole-brain average) signal. From these ROI time series, we regress out motion confounds, low-pass filter the signals below 0.15Hz, and use the 95th-percentile of the absolute amplitude to normalize the demeaned ROI time courses. To prepare the EEG data for input, we first exclude the ECG, EOG, and EMG channels (remaining 26 channels), and resample the EEG to 200 Hz to enhance computational efficiency while preserving meaningful frequency components (which are typically below 100 Hz). To predict the fMRI ROI signal, we extract EEG windows of 16 seconds before each fMRI data point. This window length is selected to encompass the peak and most of the variation in the hemodynamic response function. Additionally, this duration aligns with the maximum temporal encoding window length of the pre-trained LaBraM model Jiang et al. (2024), ensuring optimal fine-tuning. The normalization of the EEG also strictly follows Jiang et al. (2024), where EEG data values are divided by 100 so that the resulting amplitude falls primarily between -1 to 1, since a typical EEG amplitude range is from -100 $\mu V$ to 100 $\mu V$.
@@
The baseline models include state-of-the-art EEG encoding models from Yang et al. (2023) and Jiang et al. (2024), as well as two EEG-to-fMRI translation baselines Kovalev et al. (2022); Li et al. (2024). Since the original EEG-to-fMRI baselines are sequence-to-sequence models, we adapted them by adding a final projection layer to their encoder to enable sequence-to-one prediction. Additional details on baseline implementations are provided in the Appendix E.2. We employed the following two metrics for comparison: 1) Pearson correlation coefficient (R), which measures the strength and direction of the linear relationship between prediction and ground truth; 2) Mean squared error (MSE), which measures the average squared differences between prediction and ground truth across samples. The main text highlights the correlation coefficient as the primary evaluation metric, while MSE results are provided in the Appendix Table 6.
@@
For the spatiotemporal module, we initialize the model by loading the pretrained weights from LaBraM-base Jiang et al. (2024), with a token length = 200, i.e., 1 second, with no overlap. Please refer to Jiang et al. (2024) for more details of EEG pretraining. For the multi-scale spectral module, we set the smallest scale size $l_{0}=100$, i.e., 0.5 seconds without overlap. Experiments are conducted on a single RTX A5000 GPU using Python 3.9.12, Pytorch 2.0.1, and CUDA 11.7. To ensure consistency during model training, we set a fixed seed across all experiments. The batch sizes are set at 16 and 64 for intra-subject and inter-subject analyses, respectively. AdamW is utilized as our optimizer, and MSE as our training objective. The initial learning rate is set at 3e-4 with a weight decay of 0.05, and a minimal learning rate of 1e-6. For subject-specific prediction, where training and testing occur on the same scan, we split the scan in an 8:1:1 ratio for training, validation, and testing, i.e., training on the first 80% of the data and testing on the last 10%. Given that the fMRI signal exhibits significant autocorrelation, typically extending from about -10 to 10 seconds Chang et al. (2008), we implement gaps of 20 seconds between the training and validation sets, as well as between the validation and testing sets, to prevent data leakage. For the inter-subject analysis in resting-state data, we randomly divided the datasets into training/validation/testing sets by approximately 3:1:1 (18 scans : 5 scans : 6 scans). For the task data, we split the scans by 9 scans : 3 scans : 4 scans. Since data from the same individual might have shared representations, the two scans from the same individual are ensured to be in the same set (either training, validation, or testing set) to avoid possible data leakage. All models are optimized on the training set and evaluated on the test set, with the best model and hyperparameters selected based on the validation set.